The Ballentine Company expects sales for June, July, and August of $54,000, $60,000, and $50,000, respectively. Experience suggests that 35% of sales are for cash and 65% are on credit. The company collects 60% of its credit sales in the month following sale, 35% in the second month following sale, and 5% are not collected. What are the company's expected cash receipts for August from its current and past sales
Answer #1

Cash receipt for the month of August is the aggregate of cash sale of August, 60% of July’s credit sale, and 35% of June’s credit sale.

Cash receipt for the month of August

June

July

August

Credit sale

54,000 × 65% = 35,100

60,000 × 65% = 39,000

50,000 × 65% = 32,500

Cash receipt:

   Cash sale

50,000 × 35% = 17,500

   60% of July’s credit sale

39,000 × 60% = 23,400

   35% of Jun’s credit sale

35,100 × 35% = 12,285

Total

53,185

Answer: Cash receipt for the month of August is $53,185.